Solve the folloing systems of equations using the substitution method and check
3a+b=4
a-2b=6
please help

Answer by stanbon(75887)

Solve the folloing systems of equations using the substitution method and check
3a+b=4
a-2b=6
--------------
Solve the 2nd equation for "a":
a = 2b+6
---
Substitute for "a" and solve for "b":
3(2b+6) + b = 4
6b + 18 + b = 4
7b = - 14
---
b = -2
---
Since a = 2b+6, a = 2*-2 + 6 = 2
-------------------------------------
Solution: a = 2, b = -2
=================================
Cheers,
Stan H.